OpenStudy (anonymous):

Given cos theta = 8/13 where 3pi/2 < theta < 2pi find the value of cos 2 theta

OpenStudy (anonymous):

well no, because that is wrong, sorry, it is \[\cos(2x)=2\cos^2(x)-1\]

OpenStudy (anonymous):

so your answer is \[2(\frac{8}{13})^2-1\]whatever that is

OpenStudy (anonymous):

my possible answers are only -0.97 0.97 0.243 or -0.243 the sign < was actually < or =